# Predicate Invention from Pixels via Pretrained Vision-Language Models

### v1

@@ 3.1 Demonstrations
The environment’s action space $A$ is defined by a finite set of controllers $\mathcal{C}$. Each controller $C((\lambda_{1},\dots,\lambda_{v}),\Theta)\in\mathcal{C}$ has a semantically-meaningful name, as well as optional discrete typed parameters $(\lambda_{1},\ldots,\lambda_{v})$ and continuous parameters $\Theta$. For instance, a controller Pick for picking up an object might have one discrete parameter of type item and a $\Theta$ that is a placeholder for a specific grasp. An action $a\in\mathcal{A}$ in a task is a controller $C\in\mathcal{C}$ with discrete and continuous arguments fully specified: $a=C((o_{1},\ldots o_{v}),\theta)$.
Final Demonstration Structure
We provide the agent with a demonstration set $\mathcal{D}$. Each demonstration consists of: (1) a training task $T\sim\mathcal{T}$, (2) a near-optimal plan that solves the task, and (3) the state sequence of that plan with values for all initial predicate ground atoms.
Recall that a task is a tuple $\langle\mathcal{O},x_{0},g\rangle$. A plan $\pi=(a_{1},\ldots,a_{m})$ is a sequence of actions $a\in\mathcal{A}$ such that successive execution of the actions in the environment, starting from $x_{0}$, results in a final state $x_{m}$ where $g$ holds. A state sequence of a plan is the set of states $(x_{1},\ldots,x_{m+1})$ encountered as that plan is executed in an environment. An example of this input is shown in Figure 2.
Given a handful (less than $15$) of these demonstrations, the agent’s objective is to invent additional predicates and a set of operator definitions in terms of the initial and learned predicates that enables the agent to plan and efficiently solve a set of held-out tesk tasks drawn from $\mathcal{T}$. Test tasks involve more objects, different object configurations, novel compositions of goals, and longer-horizons than the training tasks solved in the demonstrations.
3.2 Operators and Planning
Predicates $\Psi$ induce an abstract state space $S_{\Psi}$ of the task’s underlying continuous state space $\mathcal{X}$. Operators help specify an abstract transition model over $S_{\Psi}$.
More specifically, given a particular state $x_{t}$, as well as predicates $\Psi$, we can compute the corresponding abstract state (denoted by $s_{t}$) by selecting all the ground atoms that are “true” in $x_{t}$. Formally, $s_{t}=\textsc{Abstract}(x_{t},\Psi)\triangleq\{\underline{\psi}:c_{\underline{\psi}}(x_{t})=\text{true},\forall\psi\in\Psi\}.$ Then, each operator $\omega$ is a tuple $\omega=\langle\overline{v},P,E^{+},E^{-},\mathcal{C}\rangle$. Here, $\overline{v}$ are typed variables representing the operator’s arguments, $P$, $E^{+}$ and $E^{-}$ are sets of predicates representing operator preconditions, add effects and delete effects respectively, and $\mathcal{C}$ is a controller associated with the operator. Note importantly that the discrete parameters of $\mathcal{C}$ must be a subset of $\overline{v}$. Specifying objects of the correct type as arguments induces a ground operator $\underline{\omega}=\langle\underline{P},\underline{E^{+}},\underline{E^{-}},\underline{\mathcal{C}}\rangle$. Ground operators define a (partial) abstract transition function: given $\underline{\omega}=\langle\underline{P},\underline{E^{+}},\underline{E^{-}},\underline{\mathcal{C}}\rangle$, if $\underline{P}\subseteq s$, then the successor abstract state $s^{\prime}$ is $(s\setminus\underline{E^{-}})\cup\underline{E^{+}}$ where $\setminus$ is set difference.
A ground operator specifies a corresponding ground controller $\underline{\mathcal{C}}$. To extract an action $a$ from this ground controller, we must specify values for the parameters $\Theta$ associated with controller $\mathcal{C}$. Following previous work (Silver et al. 2023; Kumar et al. 2023; Chitnis et al. 2022), we leverage continuous parameter samplers to do this.
Given these components, we can perform planning to solve test tasks. We assume access to a planner that takes a set of predicates, operators, and samplers, as well as a task $T=(\mathcal{O},x_{0},g)$ associated with an environment, and outputs a plan that achieves the goal at the symbolic level, if such a plan exists. We implement such a planner via off-the-shelf classical planning algorithms (Helmert 2006), with details in Appendix A.1, though we note that our predicate invention framework is compatible with other planning strategies.
3.3 Predicate Invention Problem
Given demonstrations $\mathcal{D}$ structured as described in Section 3.1, our objective is to learn additional predicates ($\Psi$), operators ($\Omega$), and samplers ($\Sigma$) that enable aggressive generalization to novel test tasks from the task distribution $\mathcal{T}$ by calling a planner with these learned components. We directly leverage an existing approach to learn neural network samplers from our demonstration data (Appendix A.3). We are mainly interested in learning predicates and operators.
Predicate Invention via Hill-Climbing
We build on the framework introduced by Silver et al. (2023) for predicate invention. Their method first generates an initial pool of candidate predicates $\Psi_{\text{pool}}$ from a grammar and then selects the subset of predicates from this pool that maximizes an objective function designed to measure planning efficiency over the demonstrations, where the selection is done via a hill-climbing optimization. Predicates subselected from the pool $\Psi_{\text{pool}}$ are added to the initial set $\Psi_{\text{init}}$ to form a final set of predicates $\Psi$. After, operators are learned in terms of the predicates $\Psi$. Intuitively, optimizing their particular objective fuction yields the predicate set (and corresponding operator set) that produces plans that most closely align with the demonstrations on the training tasks.
Importantly, this previous work makes two strong assumptions that do not hold in our setting: (1) the state space $\mathcal{X}$ is composed entirely of continuous object features (i.e., $\mathcal{X}$ only consists of $x^{\text{feat}}$), and (2) learned predicate classifiers are not noisy. We relax these assumptions and enable predicates that operate over images.
4 Inventing Predicates from Pixels
Our main interest is to invent predicates from pixels (i.e., predicates that operate directly over one or more images in our image-based state $x^{\text{img}}$). We wish to do this from a handful of structured demonstrations described in Section 3.1. These demonstrations do not possess any direct supervision for additional predicates (i.e., the demonstrations provide no indication of the number or structure of any predicates we should invent). The hill-climbing framework for predicate invention (Section 3.3) is able to learn predicates in such a setting, but requires a pool of initial candidate predicates to subselect from. To define such an initial pool, we must specify two things: what the predicates should be (i.e., what visual concepts are relevant for decision-making), and how should they be implemented (i.e., for every predicate $\psi$ in the pool, how can we define a corresponding classifier function $c_{\psi}$).
We propose to use a VLM to both generate an initial pool of predicates that might be relevant to decision-making, and implement the classifier function for each predicate in the pool. Recent work has demonstrated that VLMs are able to answer a variety of common-sense natural language questions given one or more images (Antol et al. 2015; Kim et al. 2024; Majumdar et al. 2024). The core intuition motivating our approach is that predicate generation and implementation correspond to two different types of questions. Predicate generation can be accomplished by giving the VLM access to our demonstrations, and asking “what visual concepts are changing between false and true after each action execution?”. Predicate classifier implementation can be accomplished by giving the VLM access to a particular image and asking it “is this particular ground predicate true or false in the current state?”. In the remainder of this section, we explain how exactly we implement this procedure and how we modify and leverage the hill-climbing predicate invention framework described in Section 3.3 to yield a compact subset of predicates optimized for planning. [^2] See Figure 8 for a detailed overview of our entire pipeline.
@@ 4.1 Implementing Visual Predicates with a VLM
Recall that each state $x\in\mathcal{X}$ is composed of two parts: a feature-based state $x^{\text{feat}}$ and an image-based state $x^{\text{img}}$. Visual predicates operate exclusively over $x^{\text{img}}$. Specifically, a visual predicate is a tuple of a name, a sequence of $m$ types, and a classifier function: $\psi_{\text{vis}}=\langle\text{name},(\lambda_{1},\dots,\lambda_{m}),c_{\psi}\rangle$. Here, the classifier function returns a boolean given an image-based state and object arguments $c_{\psi}:\mathcal{X}^{\text{img}}\times\mathcal{O}^{m}\to\{\text{true},\text{false}\}$.
We propose to implement the classifier function $c_{\psi}$ by querying a VLM. Specifically, we create a prompt string $\text{txt}_{\psi}$ with the predicate’s name and arguments to be filled in by the predicate’s typed object arguments $(\lambda_{1},\dots,\lambda_{m})$ (e.g. Cooked(?obj1)). [^3] We create the prompt string $\text{txt}_{\psi}$ with the predicate’s name but other choices are possible. We could ask the VLM, “Is patty1 cooked?” instead of “What is the value of Cooked(patty1)?") When the predicate is ground, we substitute the specific objects used to ground it into the prompt (e.g. Cooked(patty1)). We implement the classifier function by simply passing the prompt with this ground atom string into a VLM with some additional instruction, asking it to output either “true”, “false”, or “unknown” (which we take to be equivalent to “false”).
While this idea is straightforward, we found that several additional modifications to the text prompt $\text{txt}_{\psi}$ and images $x^{\text{img}}$ were critical to obtain accurate labels for ground atoms from the VLM. We discuss these in detail and provide examples of our prompts in Appendix A.5.
4.2 Proposing an Initial Pool of Visual Predicates
Figure 3: VLM predicate proposal example. We prompt (Appendix A.5) a VLM to propose a large pool of predicates that describe important concepts in the given demonstrations. Technically, the VLM outputs ground atoms as strings, and we parse the predicate (visualized in orange) from each ground atom. Shown here are proposals generated from demonstrations in the task distribution “Combo Burger” within the “Burger” environment (Section 5). The VLM proposes 563 predicates and we filter these into 136 “valid” predicates; the figure shows 56 of these.
Figure 4: VLM predicate labeling example. We prompt (Appendix A.5) a VLM to determine the truth values of ground atoms (each represented as a string) in the given image, labeling each as either true, false, or unknown. Shown here are a subset of the labels generated for a particular state from a task in the task distribution“Combo Burger” within the “Burger” environment (Section 5).
We generate a pool candidate of visual predicates by prompting a VLM on each demonstration $d\in\mathcal{D}$. For a demonstration $d\in\mathcal{D}$ of length $k=len(d)$, we extract the image-based state at each timestep $x^{\text{img}}_{t},t\in[0,k-1]$ and add a text heading to each image corresponding to which timestep $t$ in the demonstration it belongs to. We then pass all these images in sequence, along with the actions executed in between them ($a_{0},a_{1},\ldots,a_{k_{d}-1}$) directly to a VLM. We prompt the VLM to output a set of proposals for ground atoms based on these inputs.
Given a set of ground atoms proposed for each demonstration, we then parse this set to discover atoms that are syntactically incorrect. In particular, we remove atoms that include object names that are not part of the demonstration $d$ (for instance, the VLM might propose an atom ontop(patty1, pot1), even though there is no object named pot1). We then lift each of these atoms into a visual predicate by creating typed variables for each of the objects identified in each proposed ground atom (for instance, an atom cooked(patty1) would be lifted to cooked(?p: patty)). Finally, we remove any duplicate lifted atoms and add only unique predicates to our pool. We obtain a final pool of predicates $\Psi_{\text{pool}}^{\text{vis}}$ by running this procedure on each demonstration.
4.3 Predicate Invention via Subset Selection From Noisy Data
Given the procedure described in Section 4.2, we can generate an initial pool of visual predicates $\Psi_{\text{pool}}^{\text{vis}}$. We combine these with feature-based predicates (generated via a programmatic grammar as described in Silver et al. (2023)) to create an overall pool $\Psi_{\text{pool}}\leftarrow\Psi_{\text{pool}}^{\text{vis}}\cup\Psi_{\text{pool}}^{\text{feat}}$. We wish to run hill-climbing optimization as described in Section 3.3 to subselect a small set of relevant predicates that are optimized for efficient and effective planning over our training tasks.
Unfortunately, we cannot directly apply the particular hill-climbing optimization from previous work because our visual predicates are noisy. Visual predicates might not be consistent across states in the demonstrations due to occasional hallucination or mislabeling on the part of the VLM. This noise creates outliers in the abstract state space transition data and causes the operator learning algorithm to overfit and create a large number of unnecessary operators to model these outlier transitions. These extra operators make planning inefficient by increasing the branching factor of the planner’s search algorithm. They also cause the hill-climbing predicate subselection – which interally performs operator learning on every subset of predicates it scores – to overfit and select a large number of unnecessary predicates. These overfit predicates can appear in the definitions of all operators – not just the operators created to model the outlier transition data – because operators are learned to be chainable (add effects of an earlier operator satisfy the preconditions of a later operator). The overfit predicates and overfit operators fail to generalize outside of the training distribution. We combat this problem by modifying the operator learning approach to be more tolerant to noise  [^4] Modifications are described in detail in Appendix A.2, and by regularizing the hill-climbing optimization itself via early stopping to discourage overfitting.
@@ 5 Experiments
Our experiments are designed to answer the following questions.
@@
    How well does pix2pred generalize to novel, more complex goals when compared to an imitation approach that doesn’t use a planner?
@@
  •       Q3.
    How critical is it to invent both visual and feature-based predicates?
Figure 6: Environments. Top row: train task examples. Bottom row: test task examples. See Section 5 for descriptions of each environment.
Environments. We now describe our experimental environments and tasks, with details in Appendix A.6. We implement 4 task distributions across 2 environments. Each environment was introduced by previous work and lightly adapted for our setting.
@@
    Kitchen: A simulated robotic environment featuring a robotic arm positioned in front of a kitchen that includes a stove, a microwave, and controllable lights (Gupta et al. 2019). The task is to boil water in a kettle by turning on a specific stove burner and pushing the kettle onto it. The task has a short horizon, requiring only two high-level skills in sequence to achieve the goal. Demonstrations show the robot solving the task using a particular burner from various kettle initial positions, and at test time the robot must use a different burner on the same stove.
@@
    Burger: A simulated 2D grid world environment (based on the Robotouille environment introduced by (Wang et al. 2023)) featuring a robot that must prepare and assemble burgers from various components. Preparation includes chopping lettuce on a cutting board and cooking patties on a grill. We experiment with three task distributions within this environment, with each distribution featuring its own set of training and testing tasks: (1) Bigger Burger involves producing a burger with extra patties after being shown how to create a burger with a single patty, (2) More Burger Stacks involves producing multiple burgers after being shown how to create a single burger, and (3) Combo Burger involves producing burgers with multiple ingredients after being shown how to create burgers with a single ingredient. The tasks have a long horizon. For example, test tasks in Bigger Burger, More Burger Stacks, and Combo Burger can require a sequence of 16, 29, and 30 controllers to achieve the goal, respectively.
Figure 7: pix2pred versus baselines. Percent of evaluation tasks solved per task across both simulated environments. All results are averaged over 5 seeds. Black bars denote standard deviations.
Approaches. We evaluate our method against several ablations and baselines from the literature, including those that attempt to solve test tasks directly without learning any abstractions.
@@
    Ours: Inventing visual and feature-based predicates with pix2pred, learning operators in terms of these predicates, and planning with these operators at test time.
@@
    VLM subselect: An ablation of pix2pred where we have the VLM directly subselect a compact set of visual predicates instead of delegating subselection to the hill-climbing optimization. We still learn feature-based predicates via hill-climbing in this case.
@@
    No subselect: An ablation of pix2pred where we perform no hill-climbing, but rather select all predicates in $\Psi_{\text{pool}}$.
@@
    No invent: An ablation of pix2pred where we do not propose or subselect any additional predicates beyond the provided ones in $\Psi_{\text{init}}$.
@@
    No visual: An ablation of pix2pred where we only invent feature-based predicates.
@@
    No feature: An ablation of pix2pred where we only invent visual predicates.
@@
    VLM feat. pred: A baseline inspired by recent work from Han et al. (2024). Here, a VLM writes predicates as functions over the feature-based state instead of being able to invent its own novel visual concepts.

### v4

@@ 3.1 States, Actions, and Objects
AA is the agent’s action space defined by a finite set of skills $\mathcal{C}$ that are temporally-extended low-level robot behaviors. Each skill $C((\lambda_{1},\dots,\lambda_{v}),\Theta)\in\mathcal{C}$ has a semantically-meaningful name, a policy function, optional discrete typed object parameters $(\lambda_{1},\ldots,\lambda_{v})$, and optional continuous parameters $\Theta$. For instance, a skill Grasp for picking up an object might have one discrete parameter of type object and a 6-dimensional $\Theta$ that is a placeholder for a specific grasp. An action $a\in\mathcal{A}$ is a skill $C\in\mathcal{C}$ with discrete and continuous arguments fully specified: $a=C((o_{1},\ldots o_{v}),\theta)$. An action executes an underlying policy from the policy function (e.g. a learned grasp policy) to modify the state.
Figure 3: Domains. Top row: train task example illustrations. Bottom row: test task example illustrations.
3.2 Predicates, Symbolic World Models, and Planning
A predicate is characterized by a name, an ordered list of $m$ arguments $(\lambda_{1},\dots,\lambda_{m})$ (note that the types of the arguments are not necessarily unique). and a classifier function $c_{\psi}:S\times\Lambda^{m}\to\{\text{true},\text{false}\}$ parameterized by the arguments. A ground atom $\underline{\psi}$ consists of a predicate $\psi$ and objects $(o_{1},\dots,o_{m})$. Each ground atom induces a particular binary state classifier $c_{\underline{\psi}}:S\to\{\text{true},\text{false}\}$, where $c_{\underline{\psi}}(x)\triangleq c_{\psi}(x,(o_{1},\dots,o_{m}))$. For example, the ground atom Holding(spot, apple) would have the value “true” if the apple is being held by the spot. In this work, we distinguish between feature-based predicates that operate exclusively over the object-centric state, and visual predicates that operate on the image-based state.
A set of predicates $\Psi$ induce an abstract state space $S^{\Psi}$ of the underlying continuous state space $S$. In order to enable planning, we require an action model $\Omega$ that specifies an abstract transition model over $S^{\Psi}$. We learn such an action model in the form of symbolic PDDL-style [33, 10] operators that are each linked to a particular skill from the domain’s action space. In domains where skills have continuous parameters, each operator is associated with a generative sampler that proposes continuous parameters for the associated skill [6].
Planning. Together, the predicates and operators specify a symbolic abstract world model that can be used in conjunction with samplers for efficient hierarchical planning such as Task and Motion Planning (TAMP) [45, 11]. In this work, we choose to use a simpler planning procedure that does not require a precise simulator or transition model [28]. At a high level, our planning procedure takes in the initial state $s_{0}$ and first evaluates all the provided predicates $\Psi$ to yield an abstract state. Then, it calls a PDDL planner to achieve the goal from this abstract state. We can simply execute each action of this plan directly and check whether the goal is achieved at the end (which we do in all our simulation domains), or execute actions one-by-one and replan from resulting states as necessary (which we demonstrate in some of our real-world deployment). The precise details of our planning procedure are not necessary to understand what follows; we defer them to Appendix A.2.
3.3 Learning and Deployment
Learning Phase. We learn from a set of $n$ demonstrations $\mathcal{D}$, starting with a set of initial predicates $\Psi^{\text{init}}$. Each demonstration $d$ consists of (1) a set of objects $\mathcal{O}^{d}$, (2) a k-step trajectory $\tau^{d}=[s_{0},a_{0},s_{1},a_{1},\ldots,s_{k}]$ of states and actions, and (3) a conjunctive goal expression $g^{d}$ that holds true only in the final state $s_{k}$. Here, $g^{d}$ is a conjunction of ground atoms with predicates from $\Psi^{\text{init}}$, and objects from the object set $\mathcal{O}^{d}$. The initial predicate set $\Psi^{\text{init}}$ consists of predicates that are necessary to specify the demonstration goals (we denote this subset $\Psi^{G}$), and optionally some additional predicates that might be known ahead of time. As a concrete example, for the Cleanup domain in Figure 1, the initial predicate set consists of {WipedOffTable(?obj), InsideContainer(?obj1, ?obj2), Holding(?robot, ?obj), HandEmpty(?robot)}, the first demonstration has the goal WipedOffTable(play_table), and the second has the goal InsideContainer(green_block, cardboard_box).
Deployment Phase. Given (1) a new set of objects $\mathcal{O}^{\text{test}}$ with the same types as seen in the demonstrations, (2) a novel initial state $s_{0}$, and (3) a novel goal $g^{\text{test}}$ expressed using $\Psi^{\text{init}}$, we must output a sequence of $m$ actions (which we call a plan) that achieves some new state $s_{m}$ where $g^{\text{test}}$ holds. We refer to the tuple of objects, initial state, and goal as a task.
4 Learning Symbolic World Models
Given a demonstration set $\mathcal{D}$ (Section 3.3), we learn a PDDL-style [33, 10] symbolic world model $\mathcal{W}^{\mathcal{D}}$ by inventing additional predicates $\Psi^{\mathcal{D}}$, learning operators defined by both the newly-invented predicates and the initial predicates $\Psi^{\mathcal{D}}\cup\Psi^{\text{init}}$, and, where applicable, training generative samplers for continuous skill parameters [6]. We adapt prior work for sampler learning (Appendix A.4) and focus on inventing visual predicates and operators from demonstrations.
4.1 Proposing an Initial Pool of Visual Predicates
The first step of our approach is to generate a pool of candidate visual predicate names by prompting a VLM on each demonstration $d\in\mathcal{D}$. For a demonstration $d\in\mathcal{D}$ of length $k=len(d)$, we extract the image-based state at each timestep $s^{\text{img}}_{t},t\in[0,k-1]$ (i.e., the states before and after each skill execution). We add a text heading to each image corresponding to which timestep $t$ in the demonstration it belongs to. We then pass all these images in sequence, along with description of the actions (name and specific arguments) executed in between them ($a_{0},a_{1},\ldots,a_{k-1}$), and the descriptors of the objects in each demonstration, directly to a VLM. We prompt the VLM to output a set of proposals (in the form of natural language strings) for ground atoms based on these inputs.
Given a set of ground atoms proposed for each demonstration, we then automatically parse this set to discover atoms that are syntactically incorrect. In particular, we remove atoms that include object names that are not part of the demonstration $d$ (for instance, the VLM might propose an atom Inside(apple, room1), though there is no object named room1). We then deterministically lift each of these atoms into a visual predicate by creating typed variables for each object (for instance, an atom OnTop(apple, short_round_coffee_table) would be lifted to OnTop(?m: movable, ?t: table), where movable and table are input types specified as part of specifying the objects as defining the problem and environment). Finally, we remove any duplicate lifted atoms and add only unique predicates to our pool. We obtain a final pool of predicates $\Psi_{\text{pool}}^{\text{vis}}$ by running this procedure on each demonstration.
@@ 4.2 Implementing Visual Predicates with a VLM
We implement the classifier function $c_{\psi}$ for all visual predicates $\psi^{\text{vis}}$ as a query to a VLM. Specifically, we create a prompt string $\text{txt}_{\psi}$ with each predicate’s name and arguments (e.g. NoObjectsOnTopTable(?surface)). When the predicate is ground, we substitute the descriptors of the specific objects used to ground it into the prompt (e.g. NoObjectsOnTopTable (short_round_coffee_table)). To evaluate the predicates, we pass the prompt with these ground atom strings into a VLM with additional instructions, asking it to output either “true”, “false”, or “unknown” (which we take to be equivalent to “false”) for each predicate. If the current state is not the initial state (i.e., for $s_{t}:t>0$) of a task, we also provide the previous state’s image-based state, ground atom labeling, and executed action. We discuss our approach to prompting for atom labeling and proposal in detail and provide examples of our prompts in Appendix A.5.
4.3 Learning Symbolic World Models via Optimization
Algorithm 1 shows the pseudocode of our hill-climbing procedure for learning predicates and operators. Here, $\Psi^{\text{vis}}_{\text{pool}}$ is generated by the procedure described in Section 4.1. Where possible (such as in simulation domains where object features are available), we combine these with additional “feature-based” predicates that are generated from a programmatic grammar over the object-centric state space, as described in Silver et al. [42]), to create an overall pool $\Psi_{\text{pool}}\leftarrow\Psi_{\text{pool}}^{\text{vis}}\cup\Psi_{\text{pool}}^{\text{feat}}$. The feature-based predicates are intended to capture relations that are challenging to capture visually (e.g. relations involving robot joint angles). $\Psi_{\text{pool}}$ is passed into the hill-climbing procedure from [42] that optimizes a planning-based objective ($J(\ldots)$). Internally, it performs operator learning (OperatorLearn) and outputs operators in terms of a small subset of predicates from $\Psi_{\text{pool}}$.
Input: Demos $\mathcal{D}$, initial predicates $\Psi_{\text{init}}$
$\Psi_{\text{pool}}\leftarrow\Psi_{\text{init}}\cup\Psi_{\text{pool}}^{\text{vis}}\cup\Psi_{\text{pool}}^{\text{feat}}$;
$\Psi_{\text{selected}}\leftarrow\emptyset$, $J_{\text{prev}},J_{\text{curr}}\leftarrow\infty$;
while $J_{\text{prev}}-J_{\text{curr}}>J_{\text{thresh}}$ do
 $J_{\text{prev}}\leftarrow J_{\text{curr}}$;
 $\psi^{*}\leftarrow\arg\min_{\psi\in\Psi_{\text{pool}}}J(\mathcal{D},\text{OperatorLearn}(\mathcal{D},\Psi_{\text{selected}}\cup\{\psi\}),\Psi_{\text{selected}}\cup\{\psi\})$;
 $J_{\text{curr}}\leftarrow J(\mathcal{D},\text{OperatorLearn}(\mathcal{D},\Psi_{\text{selected}}\cup\{\psi^{*}\}),\Psi_{\text{selected}}\cup\{\psi^{*}\})$;
 $\Psi_{\text{selected}}\leftarrow\Psi_{\text{selected}}\cup\{\psi^{*}\}$, $\Psi_{\text{pool}}\leftarrow\Psi_{\text{pool}}\setminus\{\psi^{*}\}$;
return predicates $\Psi_{\text{selected}}\cup\Psi^{G}$, $\text{OperatorLearn}(\mathcal{D},\Psi_{\text{selected}})$
Algorithm 1 Model Learning
We make a few core modifications to the version of Algorithm 1 from [42]. We regularize the optimization by early-stopping the hill-climbing optimization via a hyperparameter ($J_{\text{thresh}}$) that we tune on a small validation set. Noise also tends to adversely impact operator learning: small differences in atom values before and after a particular skill is executed lead to learning many operators that are each overfit to specific cases, instead of one general operator that handles many transitions. We combat this by modifying the precondition induction procedure of operator learning to be more noise aware, as well as simply deleting operators that model less than a certain fraction of the data (controlled by a hyperparameter). See Appendix A.3 for more details.
A key advantage of our approach is that the optimization will automatically select among several closely-related and even synonymous predicates (such as NextTo vs. Touching in Figure 2) to find the predicates that are most accurately labeled and also important for planning across the training demonstrations.
@@ 5 Experiments
We design experiments to answer the following questions.
@@
    How well does pix2pred generalize to novel, more complex goals, especially when compared to an imitation approach that doesn’t use a planner?
@@
Domains. We now describe our experimental domains and tasks (illustrated in Figure 3), with details in Appendix A.6. We implement 2 real-world tasks on a Boston Dynamics Spot robot in different real-world buildings. For the purposes of extensive comparison against baselines, and to demonstrate the broad applicability of our method across sequential decision-making problems, we use 3 simulated domains that were introduced by previous work and lightly adapted for our setting.
@@
    Cleanup: A real-world Spot clears objects into a bin (sometimes dumping the bin) and wipes a tabletop. Testing tasks vary objects, backgrounds, and require more complex compositions of skills.
@@
    Juice: A real-world Spot operates a cold-press juicer by setting up cups, loading fruit, and running the machine. Testing tasks introduce new objects, backgrounds, and goals requiring demonstration composition.
@@
    Kitchen: Originally introduced by Gupta et al. [12]: a robot must boil water by placing a kettle on a stove burner and turning it on. Testing tasks involve using a different burner.
@@
    Burger: A simulated burger-making grid world [48]. We test three scenarios: Bigger Burger tasks require stacking extra patties, More Burger Stacks tasks involve assembling multiple burgers instead of one, and Combo Burger tasks demand combining multiple different ingredients at test time without seeing combining at training time. Test tasks involve sequences of up to 30 skill executions
@@
    Coffee: A simulated task [41] where a robot dispenses coffee into a jug from a coffee machine and pours coffee into cups of varying sizes. Test tasks add more cups and novel object poses.
Figure 4: pix2pred versus baselines in simulation. Percent of evaluation tasks solved across all our simulated domains. All results are averaged over 5 seeds. Black bars denote standard deviations.
Approaches. We evaluate the following ablations and baselines in all simulated domains.
@@
    Ours: Our full main method: pix2pred.
@@
    VLM subselect: Ablation where the VLM directly selects a compact set of visual predicates ($\Psi^{\text{vis}}$); feature-based predicates ($\Psi^{\text{feat}}$) are still selected via hill-climbing.
@@
    No subselect: Ablation with no hill-climbing; selects all predicates in $\Psi_{\text{pool}}$.
@@
    No invent: Ablation with no new predicates beyond $\Psi_{\text{init}}$.